OpenStudy (anonymous):

You roll 2 dice. What is the probability that the sum of the dice is odd or 1 die shows a 3? A 6 X 6 table of dice outcomes will help you to answer this question.

OpenStudy (anonymous):

Probability that one die has 3 = 2 (amount of dice) * 1/6 (probability of getting a 3/die) = 1/3 Sum is odd: One dice = odd & other = even To get odd, can use 1 & 5 (3 is taken) & to get even, can use 2, 4, & 6 Therefore, probability of sum = odd is 2/6*1/2 = 1/6 Add two probabilities together = 1/3 + 1/6 = 1/2

OpenStudy (anonymous):

the quiz said it was 23/36?